OpenStudy (anonymous):

Evaluate [r]; i.e., the units of r, in the expression r = squareroot of (s − a) (s − b) (s − c) over s , where s = 1/2(a + b + c), and a, b, and c have units of meters.

OpenStudy (jdoe0001):

so.... what are the values for "a", "b" and "c"?

OpenStudy (jdoe0001):

or do you have a triangle picture to show us? yes I know is heron's formula \(\bf s = \cfrac{a+b+c}{2}\\ \sqrt{s(s-a)(s-b)(s-c)}\)

OpenStudy (anonymous):

The question doesnt say the values

OpenStudy (anonymous):

[r] = m 2. [r] = m1/3 3. [r] = 1 m 4. [r] = 1 m2 5. [r] = m3 6. [r] = 1 m1/3 7. [r] = m1/2

OpenStudy (anonymous):

8. [r] = m2 9. [r] = 1 m1/2

OpenStudy (jdoe0001):

hmm, Heron's formula is used to get the Area of a triangle when all you have is the sides only

OpenStudy (jdoe0001):

can you take a screenshot of the material and post it here?

OpenStudy (anonymous):

yea give me a second

OpenStudy (anonymous):

What are the keys to take a snapshot I'm trying and it won't work

OpenStudy (anonymous):

can I put a link?

OpenStudy (jdoe0001):

sorry I didn't read you earlier yes, a reachable one :)

OpenStudy (jdoe0001):

keys to a screenshot press PrintScr button, near the F12 key on the keyboard go to Paint and Paste there

OpenStudy (anonymous):

I was able to get the answer by the way but thank you so much for trying :)

OpenStudy (jdoe0001):

then crop anything you want and dribble on the image anything you want to dribble on it, save it, and here, use the [Attach File] blue button :)